### v1

@@ 5 Discussion
From the previous results the most apparent growth variable affecting the h-BN crystal quality is the flux composition. The published CBF work shows that, for that technique, the choice of Fe or Fe-Cr yields relatively minor differences in crystal quality and has no significant impact on crystal size [40, 44, 45]. This is unexpected since the solubility of N in Fe alone is significantly lower than in Cr alone [66, 67] and the increased concentration of Cr is well known to increase N solubility in Fe-Cr alloys [68]. Consistent with the CBF results, the microCT data indicates that, with other parameters held constant at appropriately optimized values, both Fe and FeCr can yield relatively large regions of largely flux-free h-BN (see Fig. 2a versus Fig. 2c and Fig. 3a versus Fig. 3c).
@@
However, in contrast to the CBF results, optical characterization indicates that significantly higher quality is obtained from TSFZ growth without the use of Cr in the flux. In the PL data (Fig. 6a) the sharp reduction in the signals near $E$ = 3.9 eV and 4.1 eV for sample 1 (Fe flux) as compared to 2 and 3 (FeCrB0.2 flux) qualitatively indicates a higher purity. Furthermore, only sample 1 shows sharp phonon-assisted emission lines. Since samples 1 and 2 were grown under otherwise identical conditions, we conclude that the difference in flux composition leads to the differences in the PL spectrum. The Raman spectroscopy results show a similar comparison, with the FWHM of the two modes consistently narrower for samples grown with Fe or FeB0.2 (Fig. 5a and Fig. S6) as compared to those grown with FeCr or FeCrB0.2 under similar conditions (Fig. 5b, Fig. S7 and Fig. S8). Overall, these results indicate that including Cr in the flux, at least at the utilized composition, worsens the crystallinity of samples grown by TSFZ. This may be related to the formation of Cr nitrides — observed via PXRD in samples with trapped flux (Fig. S4) and in the volatilized material (not shown) — which disrupt the crystallization of h-BN. Again, these results are surprising since the optical properties of CBF samples grown with Fe and Fe-Cr are similar [40, 44, 45]. This points to a difference between the crystallization processes in the two growth techniques which further comparative experiments and characterization will be needed to understand.
@@
Pre-saturation of the flux with B does not appear to be essential to synthesizing large, flux-free regions of h-BN. Repeating the growth shown in Fig. 2c but substituting FeCr for FeCrB0.2 yielded a similar morphology (Fig. S1a) and proportion of h-BN by microCT (Fig. S2a). However, pre-saturation eliminates the need to dwell for long periods at the beginning of the growth in order to ensure that the molten flux has dissolved a suitable concentration of B. Pre-saturation also allows for tuning of the initial B concentration in the flux, which can then be maintained during the growth by setting appropriate translation rates for equalized B mass transfer.
@@
Based on previous CBF results with Ni-Cr and Fe-Cr fluxes [41, 44], it is clear that N2 gas participates in h-BN crystallization. This is sensible since the partial pressure of N2 gas controls the equilibrium concentration of N ($n_{N}$) in the metal flux according to Sieverts’ law ($n_{N}\sim\sqrt{P_{N2}}$) [69]. One might therefore expect a higher N2 pressure to increase h-BN crystallization and/or improve crystal quality by preventing N vacancies. At lower pressures this appears, qualitatively, to be the case. Growths with $P$ = 1 bar did not reliably produce as large h-BN regions as comparable growths at $P$ = 5 – 7 bar. The microCT data collected on still higher pressure growths — including the samples analyzed in Fig. 2b and 3b, as well as others for which the data is not shown — indicates that relatively large regions of h-BN can be obtained. However, increasing the pressure up to $P_{N2}$ = 50 bar does not eliminate the occurrence of trapped flux over the course of an extended growth, at least at the growth rate utilized (0.4 mm/h). Additionally, the Raman (Fig. 5c) and PL (Fig. 6a) data make clear that that the quality of the high-pressure grown sample is significantly worse than the lower pressure grown samples. The extremely broadened features at high energy in the PL data (sample 3 in Fig. 6a) suggest significant lattice disorder. This interpretation is consistent with the Raman data, in which the intralayer mode is substantially broadened compared to the lower pressure growths and the interlayer mode is completely absent. The broadening of the intralayer mode is most likely driven by an increased defect density within the $ab$-plane, which is further supported by the increased intensity of the low-energy PL signals relative to the lower pressure samples. The complete absence of the interlayer Raman mode indicates an extreme loss of coherence in the stacking order of the BN planes.
Although the previous discussion suggests that high-pressure is detrimental for growth quality, preliminary work by Sadovyi et al. on CBF growths at $P_{N2}$ = 1000 bar using Ni and Ni-Cr fluxes indicates otherwise [70, 71]. Those results show that high-quality h-BN can be synthesized by the CBF method at substantially higher $P_{N2}$ than what we have utilized. From this we can infer that the lower quality of our high-pressure growths was likely caused by the faster growth rate and not the higher pressure. This indicates that increasing the growth rate eventually likely leads to worse quality, with vgrowth= 0.4 mm/h as an upper bound, and suggests that high-quality cannot be maintained at a faster growth rate by increasing the growth pressure. At low-pressure ($P<$ 10 bar), only a growth rate of 0.1 mm/h has been proven to produce high-quality h-BN crystals. Sensitivity to the growth rate appears in-line with CBF results showing poorer quality with a faster cooling rate [72], which is analogous to the growth rate in TSFZ/LPG.
@@
Our results indicate that either of the utilized feed compositions are workable with the TSFZ geometry. However, there are tradeoffs. A B feed provides high zone stability due to superior wettability by the flux and allows faster vfeed at the same laser power due to the higher dissolution rate. TSFZ growth with a B feed rod also favors a proportionally larger polycrystalline h-BN seed rod size to maintain equalized B mass transfer, leading to a larger crystal. Recrystallization of the B rod (i.e. FZ growth) prior to usage in the h-BN growth may also improve the precursor purity as compared to the polycrystalline h-BN rods. However, even after recrystallization, some issues with thermal cracking remain, which can destabilize the growth. Furthermore, under certain growth conditions (Fe flux, $P$ = 7 bar, TSFZ) usage of a B crystal feed led to recrystallization of primarily polycrystalline h-BN (see Fig. S1c,d) along with the formation of numerous crystalline flakes (see Fig. S6). In contrast, a h-BN feed rod is robust against thermal cracking, although its wettability is worse, leading to more zone stability issues, and a slower dissolution rate. These zone stability issues are somewhat ameliorated by using a smaller diameter (3 mm) polycrystalline h-BN feed rod, thereby necessitating a matching smaller diameter polycrystalline h-BN seed rod to equalize the mass transfer of B and producing a smaller grown crystal.
@@
In contrast, for the LPG geometry, the B feed rod is essential. For a sample grown in the LPG geometry with a $D$ = 6 mm h-BN feed and a FeCrB0.2 flux (Fig. S1b), the initial region is largely flux-free h-BN (see microCT data in Fig. S2b), however the remainder of the growth is predominately flux. The cross-over likely occurred once the initial pre-saturation of B in the flux was recrystallized as h-BN, since little of the h-BN feed rod dissolved. In contrast, using a $D$ = 6 mm B crystal as the feed rod, under otherwise identical conditions, resulted in extended h-BN recrystalliztion (see Fig. S1f and Fig. S2f). Regardless of growth geometry, the successful growth of h-BN crystals using a B feed rod indicates that, as in CBF-growth [41, 44], sufficient N for the reaction can be supplied via the gas environment. An important implication of this result is that the isotope composition of crystals grown by this technique can be controlled. Isotope control, as has been reported previously [5, 47, 73], has important consequences for properties optimization.
@@
We now summarize the currently optimized growth conditions. Here, optimization is parameterized by the amount of trapped flux as indicated by the microCT data and the crystal quality as measured by Raman and PL spectroscopies. The configuration currently confirmed by these parameters to give optimal results is growth with the TSFZ geometry using an Fe flux and polycrystalline h-BN rods with vgrowth = 0.1 mm/h and $P$ = 5 – 7 bar ($P_{N2}=$ 4.5 – 6.3 bar with the given gas flow rates). This configuration yielded large volumes of h-BN crystal, with minimal flux inclusions. The quality of our best TSFZ-grown sample can be benchmarked against crystals grown by CBF through comparison of the PL spectra (Fig. 6b). The peaks corresponding to the phonon-assisted free exciton modes are comparable, with similar linewidths and relative intensities, indicating similar quality by this metric. Note that the shift in energy of these features for the enriched CBF sample is due to the change in atomic masses, which affects the energies of the phonons involved. On the low-energy side, the C defect signal at 4.1 eV is significantly stronger (5$\approx$ 5$\times$ 5 ×) in the natural abundance CBF sample than in sample 1 qualitatively indicating that, with the same flux, the TSFZ method gives higher purity. In contrast, the CBF sample grown with isotopically enriched B and N shows no resolvable peak at 4.1 eV. We speculate that this improvement in purity may be due to a lower C content in the isotopically enriched B precursor although it may also be related to the difference in flux composition. Notably, there is a clear difference in the overall shape of the background in this energy range for the TSFZ and CBF samples: both CBF samples show a broad feature centered at roughly 4.3 eV, while sample 1 shows a broad feature centered at roughly 3.9 eV. These features occur together in measurements of commercially available h-BN powders, where they are superimposed on the sharp point defect related peaks at 3.9 eV and 4.1 eV [20, 60, 61]. Prior work indicates that these broad features are related to complex DAP recombination mechanisms associated with native defects in h-BN (i.e. B or N vacancies/interstitials) [20, 60, 61]. That the TSFZ-grown sample shows only the 3.9 eV feature while the CBF-grown samples show only the 4.3 eV feature may indicate that the two features result from distinct DAP recombination mechanisms and that the different growth techniques each favor only one type of relevant defect.
@@
In the $E$ = 4.8 – 5.7 eV region we see that the TSFZ samples all show significantly higher PL intensity than the CBF ones, suggesting a higher density of the associated defects. However, it is unclear whether this behavior is intrinsic to the TSFZ growth technique or if it is driven by the difference in sample preparation prior to PL measurements. All of the TSFZ samples were sectioned out of larger boules and then mechanically polished prior to the PL measurement, unlike the CBF-grown samples, which provide suitable surfaces for PL measurements as-synthesized. Previous work has shown that the excitation spectrum of h-BN crystals changes markedly due to deformation induced by even minimal amounts of pressure [74]. Therefore, the cutting and polishing process may result in an increase in the number stacking faults causing a concomitant enhancement of the signal in this energy region.
@@ 6 Conclusions
The intrinsic characteristics of the TSFZ technique make it well-suited to addressing several of the challenges associated with h-BN single crystal growth, namely increasing sample volume while maintaining high purity and crystallographic perfection. Here, we have demonstrated the successful growth of h-BN crystals via this technique. As benchmarked by Raman and photoluminescence spectroscopies our currently established optimal growth procedure produces bulk h-BN crystals on par with, or superior to, existing CBF-grown crystals in terms of purity and crystallinity. There remain many avenues for optimization, including further exploration of the effects of growth rate, growth pressure and flux composition on crystal quality and flux incorporation. The effect of growth geometry should also be further explored via additional characterization of the LPG-grown samples. Lastly, it will be interesting to develop exfoliation methods and incorporate these crystals into various heterostructures to determine their impact on device performance. Overall, these results have important implications for successful implementation of the many technological and fundamental research applications of h-BN single crystals.

### v2

@@ 5 Discussion
From the previous results the most apparent growth variable affecting the h-BN crystal quality is the flux composition. The published CBF work shows that, for that technique, the choice of Fe or Fe-Cr yields relatively minor differences in crystal quality and has no significant impact on crystal size [40, 44, 45]. This is unexpected since the solubility of N in Fe alone is significantly lower than in Cr alone [65, 66] and the increased concentration of Cr is well known to increase N solubility in Fe-Cr alloys [67]. Consistent with the CBF results, the microCT data indicates that, with other parameters held constant at appropriately optimized values, both Fe (sample 1) and FeCr (samples 2 and 5) can yield relatively large regions of largely flux-free h-BN via TSFZ.
@@
However, in contrast to the CBF results with metal fluxes, optical characterization indicates that higher quality is obtained from TSFZ growth without the use of Cr in the flux. In the PL data (Fig. 6a) the sharp reduction in the signals near $E$ = 3.9 eV and 4.1 eV for sample 1 (Fe flux) as compared to 2 (FeCrB0.2 flux) qualitatively indicates a higher purity. Furthermore, sample 1 shows the clearest phonon-assisted emission lines. Since samples 1 and 2 in Fig. 6 were grown under otherwise identical conditions, we conclude that the difference in flux composition leads to the differences in the PL spectrum. The Raman spectroscopy results show a similar comparison, with the FWHM of the two modes consistently narrower for sample 1 than for sample 2. This is further confirmed by the more limited Raman data set (Fig. S9) from a different section of sample 2. This data also shows a broader FWHM for the intralayer mode than sample 1, confirming that this behavior is consistent throught the sample.
Overall, these results indicate that including Cr in the flux, under otherwise identical growth conditions, worsens the crystallinity of samples grown by TSFZ. This may be related to the formation of Cr nitrides, which were observed via PXRD in samples with trapped flux (Fig. S4b). It seems possible that the Cr nitrides sequester a portion of the dissolved N in the flux, shifting the flux chemistry and leading to more disorder in the recrystallized h-BN. Again, these results are surprising since the optical properties of CBF samples grown with Fe and Fe-Cr are similar [40, 44, 45] and inherent differences in the TSFZ and APHT CBF techniques are likely the cause. In the CBF technique the h-BN forms as a thin layer on the outside of the flux as it cools and the flux is not actively mixed. In contrast, material is continuously pulled from the flux in TSFZ and the flux undergoes significantly increased mixing. This difference between the techniques appears relevant since it likely influences the distribution of B and N in the flux, making them more uniform in TSFZ than in CBF growth. This may increase the formation rate of Cr nitrides and lead to an observable detrimental effect. Further experiments, characterization and modeling will be needed to fully understand these different behaviors. Such comparisons between the two techniques may prove informative about the details of the reaction pathways occurring within the flux.
@@
Presaturation of the flux with B does not appear to be essential to synthesizing large, flux-free regions of h-BN. Repeating the growth shown in Fig. 2b but substituting FeCr for FeCrB0.2 yielded a similar morphology (Fig. S1a) and proportion of h-BN by microCT (Fig. S2a). However, presaturation eliminates the need to dwell for long periods at the beginning of the growth in order to ensure that the molten flux has dissolved a suitable concentration of B. Presaturation also allows for tuning of the initial B concentration in the flux, which can then be maintained during the growth by setting appropriate translation rates for equalized B mass transfer.
[…]
Based on previous CBF results using pure B with Ni-Cr and Fe-Cr fluxes [41, 44] at atmospheric pressure, it is clear that N2 gas participates in h-BN crystallization. This is sensible since the partial pressure of N2 gas controls the equilibrium concentration of N ($n_{N}$) in the metal flux according to Sieverts’ law ($n_{N}\sim\sqrt{P_{N2}}$) [68]. Therefore a higher N2 pressure can be expected to increase h-BN crystallization and/or improve crystal quality by preventing N vacancies. At lower pressures this appears, qualitatively, to be the case. Growths with $P$ = 1 bar did not reliably produce samples with as large h-BN regions as comparable growths at $P$ = 5 – 7 bar with the same flux compostion, feed rod composition and $v_{grow}$
@@
The microCT data collected on TSFZ samples grown at $P$ = 50 bar — including sample 3 in Fig. 2c and 3c, as well as others for which the data is not shown — indicate that relatively large regions of h-BN can be obtained under these growth conditions, although the issue of trapped flux remains. However, the Raman (Fig. 5c) and PL (Fig. 6a) data for sample 3 make it clear that the quality is significantly worse than samples 1 and 2. Before discussing our interpretation of these data sets, it is important to note again that the $P$ = 50 bar growths were conducted at a significantly faster growth speed than the lower pressure growths ($v_{grow}$ = 0.4 mm/h and 0.1 mm/h, respectively), since 0.4 mm/h was the minimum obtainable speed on the high-pressure floating-zone furnace. As a result, we are unable to isolate which variable (higher $P$ or faster $v_{grow}$) is responsible for the reduced quality in sample 3. Nonetheless, the extremely broadened features at high energy in the PL data for this sample indicate significant lattice disorder. This interpretation is consistent with the Raman data, in which the intralayer mode is substantially broadened compared to samples 1 and 3 and the interlayer mode is completely absent. The broadening of the intralayer mode is most likely driven by an increased defect density within the $ab$-plane, which is further supported by the increased intensity of the low-energy ($E<$ 4.8 eV) PL signals. The complete absence of the interlayer Raman mode indicates an extreme loss of coherence in the stacking order of the BN planes in this sample.
While, based on our experiments alone, we are unable to provide a definitive conclusion about the source of the increased disorder in sample 3, preliminary work by Sadovyi et al. on CBF growths with Ni and Ni-Cr fluxes at signficantly higher $P$ = 1000 bar may provide some indication [69, 70]. The h-BN crystals produced at this pressure show very similar quality, as indicated by Raman spectroscopy, as crystals grown by CBF at $P$ = 1 bar from the same flux [71]. This indicates that, within the context of the CBF technique, increasing the pressure does not lead directly to higher quality. By analogy, this may be taken to imply that the lower quality of sample 3 is due to the faster $v_{grow}$ not the increased pressure. Sensitivity to the growth rate appears in-line with CBF results showing poorer quality with a faster cooling rate [71], which is roughly analogous to the growth rate in TSFZ/LPG. However, given the inherent differences between the two techniques, especially the different levels of physical mixing of the flux and the difference in crystallization processes (directional for TSFZ/LPG and surface only for CBF), this comparison is not definitive regarding the impacts of pressure and $v_{growth}$.
@@
Regarding the utilized feed rod compositions, our results indicate that either are workable with the TSFZ geometry. Furthermore, the choice of feed rod composition does not appear to impact the crystallinity of the resulting h-BN. Figure S10 shows Raman data collected on a sample (9) grown under similar conditions as sample 2, but with a B feed rod rather than a h-BN feed rod. The FWHM and $x_{0}$ of the modes are similar to what is observed for sample 2 (Fig. 5b) indicating similar quality by this metric. However, there are still tradeoffs with respect to the dynamics of the growth process itself. A B feed provides high zone stability due to superior wettability by the flux and allows faster vfeed at the same laser power due to the higher dissolution rate. TSFZ growth with a B feed rod also favors a proportionally larger polycrystalline h-BN seed rod size to maintain equalized B mass transfer, leading to a larger crystal. Recrystallization of the B rod (i.e. FZ growth) prior to usage in the h-BN growth may also improve the precursor purity as compared to the polycrystalline h-BN rods. However, even after recrystallization, some issues with thermal cracking remain, which can destabilize the growth. Furthermore, as mentioned previously, under certain conditions ($P$ = 5 – 7 bar, Fe or FeB0.2 flux) usage of a B feed in the TSFZ geometry led to recrystallization of primarily polycrystalline h-BN and growth of dendritic, crystalline flakes. We suspect that this behavior arises due to a rapid supersaturation of the flux with B and N which leads to consitutional supercooling at the growth front and an unstable interface. This is possibly supported by the observation of the spontaneous nucleation of h-BN "islands" on the flux surface, as seen in Fig. S1c and Fig. S1d.
In contrast, a h-BN feed rod is robust against thermal cracking, although its wettability is worse, leading to more zone stability issues, and a slower dissolution rate. These zone stability issues are somewhat ameliorated by using a smaller diameter (3 mm) polycrystalline h-BN feed rod, however this requires use of a matching smaller diameter polycrystalline h-BN seed rod to equalize the mass transfer of B which leads to a smaller grown crystal.
@@
For the LPG geometry, the B feed rod is essential. For sample 6 grown in the LPG geometry with a h-BN feed ($D_{L}$ = 6 mm) and a FeCrB0.2 flux (Fig. S1b), the initial region is largely flux-free h-BN (see microCT data in Fig. S2b), however the remainder of the growth is predominately flux. The cross-over likely occurred once the initial presaturation of B in the flux was recrystallized as h-BN, since little of the h-BN feed rod dissolved. In contrast, using a $D_{L}$ = 6 mm B crystal as the feed rod, under otherwise identical conditions, resulted in extended h-BN recrystallization for sample 10 (see Fig. S1f and Fig. S2f). Regardless of growth geometry, the successful growth of h-BN crystals using a B feed rod indicates that, as in CBF-growth [41, 44], sufficient N for the reaction can be supplied via the gas environment. An important implication of this result is that the isotope composition of crystals grown by this technique can be controlled. Isotope control, as has been reported previously [5, 47, 72], has important consequences for properties optimization.
We now summarize the currently optimized growth conditions. Here, optimization is parameterized by the amount of trapped flux as indicated by the microCT data and the crystal quality as measured by Raman and PL spectroscopies. The currently optimized configuration based on these parameters is growth with the TSFZ geometry using an Fe flux and polycrystalline h-BN rods with vgrowth = 0.1 mm/h and $P$ = 5 – 7 bar ($P_{N2}=$ 4.5 – 6.3 bar, with the given gas flow rates). This configuration yielded large volumes of crystalline h-BN with minimal flux inclusions and high-quality indicated by the spectroscopic characterization.
The quality of our best TSFZ-grown sample (1) can be benchmarked against crystals grown by the CBF techniques by comparison of the PL and Raman spectra. In Table 1 we show a comparison of quality metrics from the spectroscopic data for both variants of CBF growth and sample 1, as well as a dendritic flake from sample 8. The average FWHM of the Raman peaks for sample 1 is essentially the same as the reported values for a sample grown by the APHT CBF technique from an Fe flux [45], which are both quite close to the best value reported for crystals from the HPHT CBF technique. In Fig. 6b we make a direct comparison of PL spectra for sample 1 and CBF samples grown from the APHT technique. The peaks corresponding to the phonon-assisted free exciton modes are comparable, with similar linewidths and relative intensities, indicating similar quality by this metric. Note that the shift in energy of these features for the isotope enriched CBF sample is due to the change in atomic masses, which affects the energies of the phonons involved.
@@
On the low-energy side, the C defect signal at 4.1 eV is significantly stronger (5$\approx$ 5$\times$ 5 ×) in the natural isotope abundance CBF sample than in sample 1, qualitatively indicating that, with the same flux, the TSFZ method gives higher purity. In contrast, the CBF sample grown with isotopically enriched B and N shows no resolvable peak at 4.1 eV. We speculate that this improvement in purity may be due to a lower C content in the isotopically enriched B precursor although it may also be related to the difference in flux composition. Notably, there is a clear difference in the overall shape of the background in this energy range for the TSFZ and CBF samples: both CBF samples show a broad feature centered at roughly 4.3 eV, while sample 1 shows a broad feature centered at roughly 3.9 eV. These features occur together in measurements of commercially available h-BN powders, where they are superimposed on the sharp point defect related peaks at 3.9 eV and 4.1 eV [20, 60, 61]. Prior work indicates that these broad features are related to complex DAP recombination mechanisms associated with native defects in h-BN (i.e. B or N vacancies/interstitials) [20, 60, 61]. That the TSFZ-grown sample shows only the 3.9 eV feature while the CBF-grown samples show only the 4.3 eV feature may indicate that the two features result from distinct DAP recombination mechanisms and that the different growth techniques each favor only one type of the relevant defect.
@@
In the $E$ = 4.8 – 5.7 eV region we see that the TSFZ samples all show significantly higher PL intensity than the CBF ones, suggesting a higher density of the associated defects. This is reflected in the higher PL ratio shown in Table 1. If this behavior is intrinsic it would suggest that an even slower growth rate ($v_{grow}$ < 0.1 mm/h) is needed to reduce the stacking fault density. However, we note that the dendritic flake samples (measured without any cutting or polishing) show a significantly lowered PL ratio, comparable to the CBF samples. Furthermore, the TSFZ samples measured in Fig. 6 were sectioned out of larger boules and then mechanically polished prior to the PL measurement, unlike the CBF-grown samples, which provide suitable surfaces for PL measurements as-synthesized. Previous work has shown that the excitation spectrum of h-BN crystals changes markedly due to deformation induced by even minimal amounts of pressure [73]. Therefore, it is possible that the cutting and polishing process may result in an extrinsic increase in the stacking faults density causing a concomitant enhancement of the signal in this energy region. As a result, the intrinsic stacking fault density characteristic of the TSFZ technique is currently unclear and further PL measurements are in progress to address this point.
Table 1: A comparison of crystal quality metrics from Raman and PL spectroscopies. For the APHT (Fe flux) and HPHT CBF techniques, the best (lowest) values found from the literature are shown. For the TSFZ samples, results from sample 1 (Fig. 5a and Fig. 6a) and from the dendritic flakes found in sample 8 (see Fig. S1d, Fig. S2d and Fig. S5a) are shown. For the Raman data, we show the (average) FWHM of the intralayer E2g phonon mode. The variable "PL ratio" is defined as the intensity ratio PL(5.48 eV) / PL(5.77 eV), which correspond to features from stacking faults and the phonon-assisted free exciton luminescence, respectively.
Technique Raman FWHM [cm-1] PL ratio
CBF - APHT 7.6 [45] 2.4 (Fig. 6a), 1.1 [45]
CBF - HPHT 7.3 [74] 0.9 [75]
TSFZ - sample 1 7.7 20.9
TSFZ - sample 8 7.6 2.7
@@ 6 Conclusions
The intrinsic characteristics of the TSFZ technique make it well-suited to addressing several of the challenges associated with h-BN single crystal growth, namely increasing sample volume while maintaining high purity and crystallographic perfection. Here, we have shown the successful growth of h-BN crystals via this technique and thereby demonstrated its feasibility. As benchmarked by Raman and photoluminescence spectroscopies, our currently established optimal growth procedure produces bulk h-BN crystals on par with, or superior to, existing CBF-grown crystals in terms of purity and crystallinity. However, there remain many avenues for optimization of the TSFZ technique, as well as open questions. While the TSFZ samples reported here mostly show higher stacking fault densities than reported CBF samples, it remains to be determined whether this is intrinsic to the employed TSFZ growth parameters. Further explorations of the growth rate, increased growth pressure and flux composition are all desirable to understand how these variables interact to impact crystal quality and flux incorporation. Determining how changes in the rotation rate and thermal gradient affect the growth interface shape and grain nucleation and growth will also be helpful. Additionally, the effect of growth geometry should also be further explored via additional characterization of the LPG-grown samples. Lastly, it will be necessary to adapt current exfoliation methods to these crystals in order to incorporate them into heterostructures and determine their impact on device performance. Overall, these results have important implications for the successful implementation of the many technological and fundamental research applications of h-BN single crystals.